please assist i would greatly appreciate it:
Ching leaves San Francisco and travels toward Los Angeles at 50 mi/h. An hour later, Phil leaves Los Angeles and travels toward San Francisco at 60 mi/h. If the two cities are 380 miles apart, how many hours will it take for Ching to meet Phil?

Answer by stanbon(57346)

Ching leaves San Francisco and travels toward Los Angeles at 50 mi/h. An hour later, Phil leaves Los Angeles and travels toward San Francisco at 60 mi/h. If the two cities are 380 miles apart, how many hours will it take for Ching to meet Phil?
-----------
Ching DATA:
rate= 50 mph ; Time = x hr. ; Distance = 50x mi
----------
Phil DATA:
rate = 60 mph ; Time = x-1 hr ; Distance = 60(x-1)=60x-60 mi
-------------
EQUATION:
distance + distance = 380 mi.
50x+60x-60=380
110x=440
x=4 hrs.
They will meet 4 hrs after Ching leaaves home.
Cheers,
Stan H.